OpenStudy (anonymous):

I need to make sure that I am on the right track to a question. The question is, "A rectangular pool has an area of 880 ft^2. The length is 10 ft longer than the width. Find the dimensions of the pool. Solve by completing the square. Round answers to the nearest tenth of a foot." The equation is x^2 + 10x = 880...This is what I have so far: x^2 + 10x = 880 x^2 + 10x + 25 = 905 905= (x + 25)^2 √905= x + 25 Correct, right? However, the question states, that I need to find the dimensions of the pool, and I have to round to the nearest tenth of a foot, so what do you think?

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

905= (x + 25)^2 should be 905= (x + 5)^2

OpenStudy (anonymous):

other than that, looks good :)

OpenStudy (anonymous):

K, thanks! :). How would I find the dimensions to the problem....I'm kind of stuck.

OpenStudy (anonymous):

area=length*width width=x length=10+x area=x(10+x)=x^2+10x=880

OpenStudy (anonymous):

to find the width, √905= x + 5 x=√905-5 to find the length: x+10=√905+5

OpenStudy (anonymous):

So then the width would be 25.08 (rounded) and the length would be 35.08? Is that right?